# Multi-Task Label Discovery via Hierarchical Task Tokens for Partially Annotated Dense Predictions

### v1

@@ 4.1 Experimental Setup
Table 1: Comparison on NYUD-v2 under one-label and random-labels settings. Our method shows clear performance gain over three tasks, which is consistent with the visualization results.
@@
4.2 State-of-the-art Comparison
Comparison on NYUD-v2. We compare our method with li2022learning ; liu2019end on NYUD-v2 under both the one-label and random-labels settings, and the quantitative results are shown in Table 1. XTC li2022learning is the first work designed for partially annotated multi-task dense prediction. MTAN li2022learning is an attention-based MTL network designed for the fully supervised setting, and we train it with our setup. The quantitative results show that our method surpasses them by a large margin on all the metrics of the three tasks. More specifically, ours achieves $+9.63\%~{}\Delta_{MTL}$ and $+8.62\%~{}\Delta_{MTL}$ compared with li2022learning under the two partial-label settings, respectively. The qualitative comparison with the state-of-the-art method XTC li2022learning as shown in Fig. 5 can also confirm the superior performance of our method.
Similar to li2022learning , our HiTTs can also be applied on full-labels setting, which utilizes cross-task relations and task-token interactions to fertilize the multi-task learning process. We compare with some of the recent works, including multi-task interaction works like MTAN liu2019end , X-Task zamir2020robust and CCR yang2023contrastive ; and multi-task optimization strategies like Uncertainty kendall2018multi , GradNorm chen2018gradnorm , MGDA desideri2012multiple and DWA liu2019end . As shown in Table 2, our method still clearly surpasses all of the SOTA works ($13.29\%~{}\Delta_{MTL}$ overall), indicating the effective cross-task interaction brought by HiTTs. Additionally, with the aid of feature-level supervision loss $\mathcal{L}_{f}$, which is supported by global task tokens, our method can achieve $14.64\%~{}\Delta_{MTL}$ overall on the three tasks.
Table 2: Comparison on NYUD-v2 under full-labels settings. Our method achieves significantly better performance compared with SoTA multi-task learning works on all of the three tasks.
@@ 4.2 State-of-the-art Comparison
MTAN liu2019end 39.39 0.5696 28.89 0.03
X-Task zamir2020robust 38.91 0.5342 29.94 0.20
Uncertainty kendall2018multi 36.46 0.5376 27.58 0.87
GradNorm chen2018gradnorm 37.19 0.5775 28.51 -1.87
MGDA desideri2012multiple 38.65 0.5572 28.89 0.06
DWA liu2019end 36.46 0.5429 29.45 -1.83
XTC li2022learning 41.00 0.5148 28.58 4.87
XTC+ kendall2018multi 41.09 0.5090 26.78 7.58
CCR yang2023contrastive 43.09 0.4894 27.87 9.04
Ours 44.32 0.4813 25.76 13.29
Ours+$\mathcal{L}_{f}$ 45.47 0.4763 25.72 14.64
Table 3: Comparison on Cityscapes under one-label setting.
@@
Table 4: Quantitative comparison on PASCAL-Context under the one-label and random-labels setting. The ssl represents the semi-supervised learning strategy adopted in li2022learning , the CL and DL are the Contrastive Loss and Discriminator Loss in li2022learning .
     Setting      Model      Semseg.      Parsing.      Norm.      Sal.      Edge.      $\Delta_{MTL}$
     mIoU      mIoU      mErr      mIoU      odsF      (%)↑
     One-Label      STL      47.7      56.2      16.0      61.9      64.0      -
     MTL baseline      48.4      55.1      16.0      61.6      66.5      0.59
     MTL ssl li2022learning      45.0      54.0      16.9      61.7      62.4      -3.60
     MTL w.CL li2022learning      48.5      55.4      17.1      61.3      64.6      -1.33
     MTL w.DL li2022learning      48.2      56.0      17.1      61.7      64.7      -1.08
     XTC li2022learning      49.5      55.8      17.0      61.7      65.1      -0.36
     Ours      56.3      57.3      15.4      63.0      68.5      6.51
     Random-Labels      STL      60.9      55.3      14.7      64.8      66.8      -
     MTL baseline      58.4      55.3      16.0      63.9      67.8      -2.57
     MTL ssl li2022learning      59.0      55.8      15.9      64.0      66.9      -2.29
     MTL w.CL li2022learning      59.0      55.3      16.0      63.8      67.8      -2.40
     MTL w.DL li2022learning      57.9      55.2      16.2      63.4      67.4      -3.31
     XTC li2022learning      59.0      55.6      15.9      64.0      67.8      -2.10
     Ours      62.2      55.7      14.7      64.8      70.7      1.74
Comparison on Cityscapes. We also compare our results with li2022learning ; liu2019end on Cityscapes, under the one-label setting with both Semseg. and Depth. tasks. As shown in Table 3, our method achieves SOTA performance on both tasks, and significantly better performance on Depth ($15.09\%$ higher on AbS. compared with li2022learning ), resulting in an average gain of $+6.98\%$ in terms of $\Delta_{MTL}$. Additionally, it’s worth mentioning that our work is the only one achieving balanced performance gain on both tasks compared with STL. Qualitative comparisons are shown in Fig. 6.
Comparison on Pascal-Context. For the comparison on Pascal-Context, we consider both one-label and random-labels settings. As shown in Table 4, our method achieves clear improvement over other methods on the majority of tasks under both settings. The greatest enhanced task is Semseg, which has $+6.8$ and $+3.2$ mIoU on the two settings compared with li2022learning . Overall, our method is $+6.87\%$ and $+3.84\%$ higher in terms of $\Delta_{MTL}$ compared with li2022learning , the gain is more significant under one-label setting with fewer labels.
Figure 6: Comparisons with SOTA methods on Cityscapes.
Figure 7: Visualization of self-affinities heatmap (left) and PCA for distributions (right) of the fine-grained task tokens of the three tasks, with samples from NYUD-v2 testing set. Based on orthogonal embeddings, the affinities between different tokens are low and the distances between different token clustering centers are far, which indicates that more discriminative representations are learned with our method.
Table 5: Investigate the effectiveness of different components on NYUD-v2 testing set under one-label setting. Inter- and Intra-Task Learning are two key components in the Global Task Token Learning stage, and Orthogonal Embeddings are the key components in the Fine-grained Task Token Learning stage. $\mathcal{L}_{f}$ and $\mathcal{L}_{p}$ are HiTTs-based supervision signals imposing on feature spaces and predictions respectively, and $\mathcal{L}_{p}^{*}$ is naive pseudo label supervision without task tokens.
@@
HiTTs 32.48 59.61 0.5844 0.8382 30.0847 37.5827 23.9975 46.4790 58.2146 6.51
w/o. Orthogonal Embeddings 27.38 55.08 0.6049 0.8626 30.5904 38.1046 23.5233 45.7012 57.1902 2.13
w/o. Inter-Task Learning 31.26 57.99 0.5966 0.8592 30.2911 37.7714 22.7802 46.2590 58.0880 4.51
w/o. Intra-Task Learning 31.44 58.50 0.5910 0.8533 30.1432 37.6719 23.3251 46.4354 58.1508 5.23
w/o. Global Task Tokens $\boldsymbol{\theta_{i}}$ 30.03 58.21 0.5823 0.8389 30.0005 37.3750 23.4160 46.1824 58.2909 5.08
w/o. Fine-grained Task Tokens $\boldsymbol{\varphi_{i}}$ 30.53 57.18 0.5842 0.8565 30.0891 37.4465 23.2297 45.9603 58.0509 4.60
STL w. $\mathcal{L}_{p}^{*}$ 30.78 58.94 0.6693 0.9362 30.2420 37.8601 23.5830 46.4743 58.3739 3.03
MTL w. $\mathcal{L}_{p}^{*}$ 33.59 61.79 0.5882 0.8554 29.8174 36.9781 23.2875 45.8803 58.1061 6.89
HiTTs w. $\mathcal{L}_{f}$ 33.24 60.74 0.5708 0.8200 29.2227 36.9305 25.7968 48.8173 60.2608 9.75
HiTTs w. $\mathcal{L}_{p}$ 35.22 62.93 0.5613 0.8014 28.8852 36.4316 25.3873 49.1251 60.9806 11.57
HiTTs w. $\mathcal{L}_{p}+\mathcal{L}_{f}$ 35.81 63.22 0.5540 0.7939 28.5131 36.1738 26.4985 50.2357 61.8343 13.23
Table 6: Investigate the performance on labeled and unlabeled data on NYUD-v2 training set under the one-label setting. GT and pseudo represents the ground-truth and the pseudo supervision, respectively. Our method clearly shows effective learning on unlabeled training data.
@@
Table 7: Investigate the cross-task learning effect on NYUD-v2 under the few-shot setting.
Method Few-Shot-Semseg. Few-Shot-Depth. Few-Shot-Normal.
@@
Figure 8: Comparison of the task confidence map before and after refined by the fine-grained task tokens, which greatly encourage high-confidence predictions on all tasks (red color represents high-confidence areas). For noisy data like the second photo taken in a dark environment, this enhancement are more significant.
4.3 Model Analysis
Components of Hierarchical Task Tokens. As shown in Table 5, under the one-label setting on NYUD-v2, we give an ablation study of our key components. Generally, we analyze the role of global task tokens $\boldsymbol{\theta_{i}}$ and fine-grained tokens $\boldsymbol{\varphi_{i}}$, and core designs for token learning process, including the orthogonal embeddings (OE) (3.2), inter- and intra-task learning (3.1). The quantitative results clearly show that both hierarchies of tokens contribute to the multi-task performance and HiTTs boost the model performance by overall $+6.40\%~{}\Delta_{MTL}$ on all tasks compared with baseline. Without either hierarchy of tokens, the performance drops, especially on Semseg. For the learning process of HiTTs, the orthogonal embeddings (OE) are essential for generating representative fine-grained task tokens, and without OE, the performance will significantly drop ($-4.38\%~{}\Delta_{MTL}$), also especially on Semseg. which requires more discriminative category information. For the inter- and intra-task learning, both contribute to the learning process, and inter-task learning contributes more since it exploits useful representations from other tasks by cross-task feature-token interactions.
Effect of Hierarchical Feature Supervision and Label Discovery. We analyze the contributions from two types of pseudo supervision losses ($\mathcal{L}_{p}$ for pseudo label loss and $\mathcal{L}_{f}$ for feature supervision loss) on different hierarchies. As shown in Table 5, both methods boost multi-task performance: $+9.64\%~{}\Delta_{MTL}$ for $\mathcal{L}_{f}$ and $+11.46\%~{}\Delta_{MTL}$ for $\mathcal{L}_{p}$ compared with MTL baseline. $\mathcal{L}_{p}$ contributes more since fine-grained task tokens contain more specific and discriminative task information and are directly involved in the formation process of task predictions. The combination of both methods achieves better performance ($+13.12\%~{}\Delta_{MTL}$) than applying them separately, which validates the importance of consistently discovering supervision signals in both hierarchies. Compared with the naive pseudo-label supervision without the utilization of task tokens (denoted as $\mathcal{L}_{p}^{*}$), our HiTTs-based pseudo-label discovery is much better. Since MTL baseline shares an image backbone that learns stronger representations on all tasks, the MTL produces pseudo labels with better quality and surpasses STL a lot in performance ($+3.86\%~{}\Delta_{MTL}$). Our HiTTs-based supervision performs i) consistent label discovery in both feature and prediction space; ii) effective cross-task feature-token interactions, which furthermore enhance the quality of pseudo labels, and bring extra $+4.68\%~{}\Delta_{MTL}$ overall.
Learning Effect on Unlabeled Data. We also study the performance of our method on the labeled and unlabeled data separately on NYUD-v2 training set under one-label setting. As shown in Table 6, for data without labels, model with HiTTs generalizes better on them, especially on Depth. and Normal, and imposing hierarchical supervision can significantly boost the performance on unlabeled data.
Effect of Cross-Task Interactions. To further show the effect of cross-task learning brought by HiTTs, we develop new few-shot settings, under which one task has only a few labels while other tasks are fully labeled. We apply this setting respectively on the three tasks of NYUD-v2, namely few-shot-semseg, few-shot-depth and few-shot-normal. For each few-shot task, we have 10 shots for the model to learn. As shown in Table 7, we only show the performance of the few-shot tasks in the table, and due to the lack of label supervision, the STL performs poorly on each few-shot task: $5.80$ mIoU on Semseg, $0.9633$ AbS on Depth, and $47.5281$ mErr on Normal. Benefiting from the sharing backbone, MTL baseline performs much better, since the backbone can be fully supervised on the other two tasks, and gain stronger representations from other tasks. With the aid of HiTTs, the multi-task model can achieve an extra performance gain, since the cross-task interactions brought by intra-task learning can fertilize the unlabeled tasks in the decoding stage, which introduces more task-relevant information and discriminative representations to task features without label supervision. The gains brought by HiTTs are $+7.76\%$ on Semseg, $+14.86\%$ on Depth, and $+0.49\%$ on Normal respectively. Additionally, if we add the pseudo supervision signals to aid the learning process, the performance will be further improved: $+19.82\%$ on Semseg, $+19.63\%$ on Depth, and $+3.85\%$ on Normal compared with MTL baseline.
Figure 9: Comparison of the training and testing performance on each task with and without hierarchical Pseudo Supervision (P). The model trained with pseudo supervision converges faster on both train and test splits, and gains better performance.
Table 8: Comparison of computational cost on NYUD-v2. All of the methods use the same SegNet badrinarayanan2017segnet backbone.
Model Params (M) FLOPs (G) Memory (G)
MTL baseline 25.06 68.81 9.36
MTAN liu2019end 44.23 217.26 43.46
XTC li2022learning 39.84 223.16 31.42
Ours 25.32 83.14 21.19
Figure 10: Quantitative analysis of the quality of pseudo labels generated by the fine-grained task tokens on Cityscapes.
Figure 11: Comparisons of task score maps produced by global task tokens and fine-grained task tokens. The upper part is the visualization of samples on NYUD-v2 while the lower part is on Cityscapes. The score maps indicate the response of task features to task tokens, those produced by global task tokens are relatively noisy and monotonous, while those generated by fine-grained task tokens have finer granularity and less noise, which shows the hierarchy of the HiTTs learning process.
Visualization Results. We visualize: i) The distributions of fine-grained task tokens on each task in Fig. 7, with the aid of OE, the self-correlation map of tokens will be more diagonal, and the distributions after PCA have better clusters in 3-dimensional feature space. ii) The contribution of fine-grained tokens $\boldsymbol{\varphi_{i}}$ to the confidence map, we visualize the confidence map $\boldsymbol{M_{i}^{p}}$ in Fig. 8, compared with the preliminary confidence masks without fin-grained tokens $\boldsymbol{\varphi_{i}}$, the final masks have higher confidence scores after the refinement. iii) The learning curves of metrics on every task in Fig. 9, both training and testing performance are boosted consistently on all tasks with the discovered pseudo supervision (P) on both hierarchies. iv) The visual quality of pseudo labels generated by fine-grained tokens $\boldsymbol{\varphi_{i}}$ on Cityscapes are show in Fig. 10. v) The visualization of score maps produced by global task tokens and fine-grained task tokens respectively on each task. As shown in Fig. 11, we conduct visualization on both NYUD-v2 and Cityscapes datasets. The score maps indicate the response of task features to task tokens, and the response patterns of feature maps on different tasks are very different, e.g. Semseg. features highlight areas with distinguish semantics, Depth. features focus on areas with a certain depth range and Normal. features focus on surfaces with the same orientation. Comparing the score maps produced by tokens from different hierarchies, we find that score maps produced by global task tokens are relatively rough and noisy, while those generated by fine-grained task tokens have finer granularity and less noise, which shows the hierarchy of the HiTTs learning process. Also, we observed that the high-lighted areas of global task tokens are monotonous, while fine-grained task tokens can highlight more various details. This phenomenon is clearly observed in Cityscapes, since the ground truths of this dataset follow the long-tail distribution, thus the global task-tokens tend to learn the category with more pixel samples, and consequently always highlight the road area as shown in Fig. 11. Oppositely, the fine-grained tokens can give attention to more details, including the vehicles and pedestrians with fewer pixel samples. Thus, it is necessary to design a hierarchical structure for tokens to learn representations with different granularity.
Analysis on Computational Cost. To prove that our HiTTs are highly effective but efficient as well, in Table 8, we analyze the number of trainable parameters, average computational costs for each input image, and the total GPU memory usage during the training phase. MTAN liu2019end has a complex encoding-decoding cross-task interactive structure based on attention, which causes larger model parameters and computational overhead in both the training and inference stages. li2022learning  proposes a cross-task consistency module to regularize the training process, which causes extra parameters ($+14.78M$) and FLOPs ($+154.35G$). Our method is much less expensive compared with li2022learning , which only requires $+0.26M$ more parameters and $+14.33G$ FLOPs. The total memory cost of the baseline is $9.36G$. The best-performing method XTC li2022learning increases the memory upon the baseline to $31.42G$, while ours upon the same baseline is only $21.19G$, confirming the efficiency of our model.
5 Conclusion
Supplementary File for ‘Multi-Task Label Discovery via Hierarchical Task Tokens for Partially Annotated Dense Predictions’
Jingdong Zhang *, Hanrong Ye, Xin Li, Wenping Wang, Dan Xu
In this supplementary document, we present: (i) more details about the multi-scale global task token learning and task loss formation, (ii) more comprehensive explanation on experimental implementations, and (iii) more quantitative and qualitative experimental results.
Figure 12: Illustrations of (a) Single-scale Global Task Token Learning and (b) Multi-scale Global Task Token Learning. For the multi-scale features produced by the shared backbone, we use linear layers to produce corresponding task tokens for each scale. Then, in each scale, we query task features and conduct inter-task learning to gain multi-task multi-scale representations. The multi-scale features and tokens are fused before intra-task learning to transfer cross-scale information.
Appendix A Model Details
Appendix A.1 Discrete Quantization and Task Losses
In Sec. 3 of the body part, we have discussed how to learn HiTTs. For continuous regression tasks, such as Depth and Normal, we first need to perform a discrete quantization of the label space to provide discriminative supervision for tokens. The goal for quantization is to assign meaningful category bins to each fine-grained token for classification. As analyzed in bruggemann2021exploring , this quantization only changes the way of predicting regression task, but does not contribute to the learning performance. For depth estimation, we follow the setting in bruggemann2021exploring ; li2018monocular , and divide the range of depth values into several logarithmic bins. Our predicted task logits score $\boldsymbol{G_{i}^{\prime}}$ is used to calculate the soft-weighted sum with each bin and produce final task predictions accordingly. For surface normal estimation, we follow bruggemann2021exploring ; zeisl2014discriminatively and use K-means to learn several unit normal vectors, which serve as clustering centers, and they are also used to generate predictions with $\boldsymbol{G_{i}^{\prime}}$. This process can be expressed as:
$\boldsymbol{\hat{Y_{i}}}=\operatorname{Sum}\left(\boldsymbol{c_{i}}^{\top}\times\operatorname{Softmax}(\boldsymbol{G_{i}^{\prime}})\right),$ (10)
where $\boldsymbol{c_{i}}\in\mathbb{R}^{C_{p}}$ represents the center of each bin. In our experiments, the numbers of depth bins and normal cluster centers on NYUD-v2 are $30$ and $20$, respectively. For Cityscapes, we consider $100$ depth bins as the Cityscapes dataset is captured from outdoor scenarios and have more significant changes in depth. For PASCAL-Context, we select $40$ different unit verctors uniformly from the space to serve as normal cluster centers.
$\operatorname{L_{i}^{reg}}\left(\boldsymbol{\hat{Y_{i}}},{\boldsymbol{Y_{i}}}\right),$ (11)
$\operatorname{L_{i}^{cls}}\left(\boldsymbol{G_{i}},{\boldsymbol{Y_{i}^{oh}}}\right)=-\boldsymbol{Y_{i}^{oh}}\cdot\operatorname{log}\left(\operatorname{Softmax}(\boldsymbol{G_{i}^{\prime}})\right),$ (12)
and the overall mixture loss for each task can be written as:
$\operatorname{L_{i}}(\cdot)=\lambda_{i}\operatorname{L_{i}^{reg}}\left(\boldsymbol{\hat{Y_{i}}},{\boldsymbol{Y_{i}}}\right)+\operatorname{L_{i}^{cls}}\left(\boldsymbol{G_{i}^{\prime}},{\boldsymbol{Y_{i}^{oh}}}\right),$ (13)
Appendix A.2 Multi-scale Global Task Token Learning
For the implementations on NYUD-v2 and Cityscapes, we use the SegNet badrinarayanan2017segnet to serve as the shared backbone, which produces single-scale shared features for per-task decoding. However, on PASCAL-Context, we use ResNet-18 he2016deep as the shared backbone, which can produce multi-scale shared features for decoding. Directly fusing them ignores the various task information maintained in different scales. Thus, we propose to learn the global task tokens on different scales in order to learn more comprehensive task relations.
$\boldsymbol{\theta_{i}}^{(j)}=\boldsymbol{W^{s\to m}_{i}}^{(j)}\times\boldsymbol{\theta_{i}},\quad j=0,1,2,3,$ (14)
[…]
$\boldsymbol{\Theta}^{(j)}$ $=\left[\boldsymbol{\theta_{1}}^{(j)};\boldsymbol{\theta_{2}}^{(j)};\cdots;\boldsymbol{\theta_{T}}^{(j)}\right]^{\top}.$ (16)
[…]
$\boldsymbol{F_{i}^{\prime}}$ $=\operatorname{Conv_{i}^{1\times 1}}\left(\left[\boldsymbol{F_{i}^{\prime}}^{(0)};\boldsymbol{F_{i}^{\prime}}^{(1)};\boldsymbol{F_{i}^{\prime}}^{(2)};\boldsymbol{F_{i}^{\prime}}^{(3)}\right]\right),$ (17)
$\boldsymbol{\theta_{i}^{\prime}}$ $=\sum_{j=0}^{3}\left(\boldsymbol{W^{m\to s}_{i}}^{(j)}\times\boldsymbol{\theta_{i}^{\prime}}^{(j)}\right).$ (18)
Figure 13: Visualization of self-affinities heatmap (left) and PCA for distributions (right) of fine-grained task tokens of the two tasks on Cityscapes validation set. With orthogonal embeddings, the affinities between different tokens are low and the clustering of token distributions on each category is better, which has a similar phenomenon to the visualization in Fig. 7 of the body part.
@@ Appendix A.2 Multi-scale Global Task Token Learning
Table 9: Comparison of HiTTs with Single-scale (SS) and Multi-scale (MS) Global Task Token Learning on PASCAL-Context under the one-label and random-labels setting.
@@
Figure 14: Comparisons of task score maps produced by global task tokens and fine-grained task tokens. The upper part is the visualization of samples on NYUD-v2 while the lower part is on Cityscapes. The score maps indicate the response of task features to task tokens, those produced by global task tokens are relatively noisy and monotonous, while those generated by fine-grained task tokens have finer granularity and less noise, which shows the hierarchy of the HiTTs learning process.
Figure 15: Comparisons with SOTA works on NYUD-v2 (upper part) and Cityscapes (lower part).
Appendix B Implementation Details
Model Setting. We use SegNet badrinarayanan2017segnet as the image backbone for our experiments on NYUD-v2 and Cityscapes; and we use ResNet-18 as the image backbone, Atrous Spatial Pyramid Pooling (ASPP) chen2018encoder as the task-specific decoding heads. For the threshold $\tau_{i}$ which selects binary mask $\boldsymbol{M_{i}^{p}}$, we use {Semseg: 0.9, Depth: 0.45, Normal: 0.6 } for NYUD-v2 one-label setting, {Semseg: 0.9, Depth: 0.7, Normal: 0.7} for NYUD-v2 random-labels setting. {Semseg: 0.9, Depth: 0.5} for Cityscapes one-label setting, and {Semseg: 0.9, Parsing: 0.85, Normal: 0.7, Sal: 0.7, Edge: 0.9} for PASCAL-Context one-label and random-labels settings.
Appendix C More experimental results
Appendix C.1 More Quantitative Results

### v2

@@ B.3. Model Setting
Table 6. Comparison on NYUD-v2 under one-label and random-labels settings. Our method shows clear performance gain over three tasks, which is consistent with the visualization results.
@@
Table 7. Comparison on NYUD-v2 under full-labels settings. Our method achieves significantly better performance compared with SoTA multi-task learning works on all of the three tasks.
@@
MTAN (Liu et al., 2019) 39.39 0.5696 28.89 0.03
X-Task (Zamir et al., 2020) 38.91 0.5342 29.94 0.20
Uncertainty (Kendall et al., 2018) 36.46 0.5376 27.58 0.87
GradNorm (Chen et al., 2018a) 37.19 0.5775 28.51 -1.87
MGDA (Désidéri, 2012) 38.65 0.5572 28.89 0.06
DWA (Liu et al., 2019) 36.46 0.5429 29.45 -1.83
XTC (Li et al., 2022) 41.00 0.5148 28.58 4.87
XTC+Uncertainty (Kendall et al., 2018) 41.09 0.5090 26.78 7.58
CCR (Yang et al., 2023b) 43.09 0.4894 27.87 9.04
Ours (HiTTs) 44.32 0.4813 25.76 13.29
[…]
B.4. Evaluation Metrics
We have briefly introduced our evaluation metrics for multiple dense prediction tasks in Sec.4.1. We provide a more detailed description as follows: (i) mIoU: mean intersection over union; (ii) pAcc: per-pixel accuracy; (iii) AbS or AbR: absolute error or absolute-relative error; (iv) rmse: root mean square error (for Normal. we calculate the mean square error of the predicted angles with the ground-truths); (v) mErr: mean of angle error; (vi) odsF: optimal dataset F-measure (Martin et al., 2004); (vii) threshold: for surface normal estimation, we calculate the proportion of pixels with angle error smaller than three thresholds $\eta\in\left\{11.25^{\circ},22.50^{\circ},30^{\circ}\right\}$.
Additionally, to better evaluate the proposed method, we also consider using $\Delta_{MTL}$ proposed by (Vandenhende et al., 2021) to evaluate the overall improvement of the multi-task performances of all the tasks, which is defined as:
(16) $\Delta_{MTL}=\sum_{N}(-1)^{l_{t}}\left(M_{m,t}-M_{s,t}\right)/M_{s,t}$
Table 8. Comparison on Cityscapes under one-label setting.
@@ B.4. Evaluation Metrics
Table 9. Investigate the effectiveness of different components on NYUD-v2 testing set under one-label setting.
@@
+HiTTs w/o. OE 27.38 55.08 0.6049 0.8626 30.5904 38.1046 23.5233 45.7012 57.1902 2.13
+HiTTs w/o. Inter 31.26 57.99 0.5966 0.8592 30.2911 37.7714 22.7802 46.2590 58.0880 4.51
+HiTTs w/o. Intra 31.44 58.50 0.5910 0.8533 30.1432 37.6719 23.3251 46.4354 58.1508 5.23
+HiTTs w/o. $\boldsymbol{\theta_{i}}$ 30.03 58.21 0.5823 0.8389 30.0005 37.3750 23.4160 46.1824 58.2909 5.08
+HiTTs w/o. $\boldsymbol{\varphi_{i}}$ 30.53 57.18 0.5842 0.8565 30.0891 37.4465 23.2297 45.9603 58.0509 4.60
+HiTTs 32.48 59.61 0.5844 0.8382 30.0847 37.5827 23.9975 46.4790 58.2146 6.51
STL w. $\mathcal{L}_{p}$ 30.78 58.94 0.6693 0.9362 30.2420 37.8601 23.5830 46.4743 58.3739 3.03
MTL w. $\mathcal{L}_{p}$ 33.59 61.79 0.5882 0.8554 29.8174 36.9781 23.2875 45.8803 58.1061 6.89
+HiTTs w. $\mathcal{L}_{f}$ 33.24 60.74 0.5708 0.8200 29.2227 36.9305 25.7968 48.8173 60.2608 9.75
+HiTTs w. $\mathcal{L}_{p}$ 35.22 62.93 0.5613 0.8014 28.8852 36.4316 25.3873 49.1251 60.9806 11.57
+HiTTs w. $\mathcal{L}_{p}+\mathcal{L}_{f}$ 35.81 63.22 0.5540 0.7939 28.5131 36.1738 26.4985 50.2357 61.8343 13.23
Table 10. Investigate the performance on labeled and unlabeled data of NYUD-v2 training set under one-label setting.
@@
Table 11. Investigate the cross-task learning effect on NYUD-v2 under the few-shot setting.
  Method   Few-Shot-Semseg   Few-Shot-Depth   Few-Shot-Normal
@@
Table 12. Comparison of HiTTs with Single-scale (SS) and Multi-scale (MS) Global Task Token Learning on PASCAL-Context under the one-label and random-labels setting.
@@
B.5. More Quantitative Results
B.5.1. State-of-the-art Comparison
Comparison on NYUD-v2. We compare our method with (Li et al., 2022; Liu et al., 2019) on NYUD-v2 under both the one-label and random-labels settings, and the quantitative results are shown in Table 6. XTC (Li et al., 2022) is the first work designed for partially annotated multi-task dense prediction. MTAN (Li et al., 2022) is an attention-based MTL network designed for the fully supervised setting, and we train it with our setup. The quantitative results show that our method surpasses them by a large margin on all the metrics of the three tasks. More specifically, ours achieves $+9.63\%~\Delta_{MTL}$ and $+8.62\%~\Delta_{MTL}$ compared with (Li et al., 2022) under the two partial-label settings, respectively.
Our HiTTs can also be applied on full-labels setting, which utilizes cross-task relations and task-token interactions to fertilize the multi-task learning process. We compare with some of the recent works, including multi-task interaction works like MTAN (Liu et al., 2019), X-Task (Zamir et al., 2020) and CCR (Yang et al., 2023b); and multi-task loss weighting strategies like Uncertainty (Kendall et al., 2018), GradNorm (Chen et al., 2018a), MGDA (Désidéri, 2012) and DWA (Liu et al., 2019). As shown in Table 7, our method still clearly surpasses all of the SOTA works ($13.29\%~\Delta_{MTL}$ overall), indicating the effective cross-task interaction brought by HiTTs. Additionally, with the aid of feature-level supervision loss $\mathcal{L}_{f}$, which is supported by global task tokens, our method can achieve $14.64\%~\Delta_{MTL}$ overall on the three tasks.
[…]
B.5.2. Model Analysis
Effect of Hierarchical Task Tokens. As shown in Table 9, under the one-label setting on NYUD-v2, we give a more-detailed ablation study on the key components of hierarchical task tokens (HiTTs), including both hierarchies of the task tokens, and the inter-task (Inter) and intra-task (Intra) learning of global task tokens, and orthogonal embeddings (OE) of fine-grained task tokens, and two types of pseudo supervision losses ($\mathcal{L}_{p}$ for pseudo label loss and $\mathcal{L}_{f}$ for feature supervision loss) on different hierarchy. The quantitative results clearly show that every component contributes to the multi-task performance on all the metrics and on all the tasks over the MTL baseline. HiTTs boost the model performance by conducting cross-task interaction and encouraging high-confidence predictions, and leads to an overall $+9.64\%~\Delta_{MTL}$ on all tasks compared with baseline. However, for the learning process of HiTTs, the orthogonal embeddings (OE) are essential for generating representative fine-grained task tokens, and without OE, the performance will significantly drop ($-4.38\%~\Delta_{MTL}$), especially on Semseg. which requires more discriminative category information. The inter-task and intra-task learning processes are also important since without either of them, the learning of cross-task relations and task representations will be affected, resulting in $-2.00\%~\Delta_{MTL}$ and $-1.28\%~\Delta_{MTL}$ performance after removing them respectively from the learning process of HiTTs.
Effect of Hierarchical Feature Supervision and Label Discovery. We analyze the contributions from both feature-level and prediction-level supervision with all of the metrics, as shown in Table 9, both methods boost multi-task performance, and $\mathcal{L}_{p}$ contributes more since fine-grained task tokens contain more specific and discriminative task information and are directly involved in the formation process of task predictions. The combination of both methods achieves better performance than applying them separately, which validates the importance of consistently discovering supervision signals in both hierarchies. Comparing $\mathcal{L}_{p}$ imposed on different models, including STL, MTL baselines, and our HiTTs, our token-based pseudo-label discovery is much better. Since MTL shares a backbone that learns stronger representations on all tasks, MTL produces pseudo labels with better quality and surpasses STL a lot in performance ($+3.86\%~\Delta_{MTL}$). Our HiTTs perform i) consistent label discovery in both feature and prediction space; ii) effective cross-task feature-token interactions, which furthermore enhance the quality of pseudo labels, and bring extra $+4.68\%~\Delta_{MTL}$ overall.
We also study the performance of our method on the labeled and unlabeled data separately on NYUD-v2 training set under the one-label setting. As shown in Table 10, for data without labels, the model with HiTTs generalizes better on them, especially on Depth. and Normal, and adding hierarchical supervision will more significantly boost the performance of unlabeled data.
[…]